OpenStudy (anonymous):

reduce the rational expression to its lowest terms. 2x^2+4x-126/3x^2-243

OpenStudy (gw2011):

2x^2+4x-126 = 2(x^2+2x-63) 3x^2-243 = 3(x^2-81) (2)(x^2+2x-63)/(3)(x^2-81) (2)(x+9)(x-7)/(3)(x+9)(x-9) = (2)(x-7)/(3)(x-9)